40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 25 10:48:13 2008, Koji, Update, General, Abs. Len. Meas. ~ Realignment / beating / PLL trial 
    Reply  Mon Jul 28 15:54:04 2008, Koji, Update, General, Abs. Len. Meas. ~ More on the beat / the PLL setup PLL_setup.pngbeat_at_8MHz.jpg
       Reply  Thu Jul 31 11:19:30 2008, Koji, Update, General, Abs. Len. Meas. ~ Resonance search trial 
          Reply  Sat Aug 2 12:53:43 2008, Koji, Update, General, Abs. Len. Meas. ~ New PLL at the PSL table yarm_dc.png
             Reply  Mon Aug 4 00:37:58 2008, Koji, Update, General, Abs. Len. Meas. ~ RF PD at the Y end / Manual frequency scan YEND_LAYOUT.png4th_FSR1.png4th_FSR2.png
                Reply  Mon Aug 4 05:23:57 2008, Koji, Update, General, Abs. Len. Meas. ~ Measurement for Y-arm completed 
                Reply  Mon Aug 4 21:48:24 2008, Koji, Update, General, Abs. Len. Meas. ~ Scan for TEM00/01/10 TEM00.pngTEM01.pngknife_edge.png
                   Reply  Mon Aug 4 22:31:10 2008, Koji, Update, General, Abs. Len. Meas. ~ Simple Test for TEM01/10 split TEM01_10.png
                   Reply  Tue Aug 5 00:05:57 2008, Koji, Update, General, Abs. Len. Meas. ~ IFR2023A calibration 
                      Reply  Wed Aug 6 11:10:34 2008, Koji, Update, General, Abs. Len. Meas. ~ analysis of the TEM00 scan TEM00fit.pngTEM00FSRfit.png
                         Reply  Wed Aug 6 11:43:52 2008, Koji, Update, General, Abs. Len. Meas. ~ analysis of the TEM01 scan TEM01fit.png
                            Reply  Thu Aug 14 21:52:51 2008, Koji, Summary, General, Abs. Len. Meas. ~ summary of my Summer mode_spacing_measurement_080816_v2.pdf
Message ID: 738     Entry time: Fri Jul 25 10:48:13 2008     Reply to this: 748
Author: Koji 
Type: Update 
Category: General 
Subject: Abs. Len. Meas. ~ Realignment / beating / PLL trial 
Alberto and Koji

o We worked for the abs length measurement setup on Thursday night.
o At the last of the work Koji left the 40m lab at 2AM. "Last autoalignment" was restored. The flipper for the 
inj beam was down. The shutter for the NPRO was closed.

----
o The alignment of the injection beam (NPRO) was re-adjusted.
o The laser crystal temp (LT) of the NPRO was scanned.
o After a long struggle the beat was found at about LT=61deg(!). I think this is almost at the maximum temp 
for the NPRO. Note that this is not the diode temp, and therefore it will not damage the laser as far as the 
TEC for the crystal works.

o Only the X arm was aligned.
o The alignment of the injection beam was adjusted such that the beating amplitude got maximum.
o At the faraday of the NPRO, we had 2.4V_DC and 1.8V_DC with and without the inj beam, respectively. The 
beating amplitude was 200mVpp (at around 2.4V).
o With a simple calculation, the mode overlapping of tghe injection beam is only 0.0023. Ahhh. It is too weak. 
In the modematching or something must be wrong. 
o The position of the mode matching lens was tweaked a little. It did not help to increase the beat ampitude. 
Even worse. (The lens was restored and the values above was obatined with the latest setting.) 

o Then tried to build a PLL. It locks easily. 
  - Put the beat signal into the mixer RF input.
  - Connect 10dBm @1MHz-10MHz from the marconi oscillator to the LO input. The supposed nominal LO level was 
not checked so far. Just used 10dBm.
  - The IF output was connected to an SR560 with 10Hz LPF (6dB/oct) with G=500 or so.We don't need to care 
about the sign.
  - The output of the SR560 was connected to FAST PZT input of the NPRO.
o The problem was that there was strong intermodulations because of 33MHz. No LPFwas used before the mixer. 
Because of this spourious modulations, the PLL servo locks at the local zero crossings. These will be solved 
next time. 

o Eventually left the 40m lab at 2AM. "Last autoalignment" was restored. The flipper for the inj beam was 
down. The shutter for the NPRO was closed.
ELOG V3.1.3-